Determine the equation of the line through (0,9) that is parallel to the line
3x + 5y =15.

Answer by tutorcecilia(2152)

Re-write the first equation in slope-intercept format (y=mx+b):
3x + 5y =15
y = -3/5x -15/5
y = -3/5x + 3
The slope (m) = (-3/5)
.
Since parallel lines have the same slope, the slope of the second line will also equal (-3/5).
.
Use the slope-intercept formula to find the equation of the second line:
y = mx + b [Plug-in the values (0, 9) and the slope (-3/5).
9 = (-3/5)(0) + b [Solve for b]
9 = b.
So, the equation of the second line is:
y = (-3/5)x + 9